OpenStudy (anonymous):

(24xy)^(-5/8) write with a positive exponents

OpenStudy (anonymous):

1 ------------ (24xy)^ (5/8)

OpenStudy (anonymous):

\[\frac{1}{24xy^\frac{5}{8}}\]

OpenStudy (anonymous):

sheg u missed the bracket !!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yeah typo error

OpenStudy (anonymous):

oh haha ,I did not think it was that simple thanks

OpenStudy (anonymous):

everything in this world is easy u just need to have that pair of eyes to look at the problems

OpenStudy (anonymous):

just remember the law 1 a^(-p/q) = -------- a^(p/q)